# Inductive Gradient Adjustment for Spectral bias in implicit neural representations

### v1

@@ 6.3 Learning 5D neural radiance fields
Learning neural radiance fields for novel view synthesis, i.e., NeRF, is the main application of INRs (Mildenhall et al., 2021; Saragadam et al., 2023; Xie et al., 2023; Shi et al., 2024a). The main process of NeRF is to build a neural radiance field from a 5D coordinate space to RGB space by an MLP $f(\mathbf{x},\Theta)$.
Table 3: Average metrics of 5D neural radiance fields by different methods. Detailed settings can be found in Sec. 6.3. Per-scene results are provided in our Appendix F.
@@
Specifically, given a ray $i$ from the camera into the scene, the MLP takes the 5D coordinates (3D spatial locations and 2D view directions) of $N$ points along the ray as inputs and outputs the corresponding color $\mathbf{c}$ and volume density $\sigma$. Then $\mathbf{c}$ and $\sigma$ are combined using numerical volume rendering to obtain the final pixel color of the ray $i$. We set each ray as one group. For each group, we sample the point with the maximum integral weight. We apply our method to the original NeRF (Mildenhall et al., 2021). The “NeRF-PyTorch” codebase (Yen-Chen, 2020) is used, and we follow its default settings for all methods. More details can be found in Appendix F.
Table 3 lists average metrics of four methods on the down-scaled Blender dataset (Mildenhall et al., 2021). Our IGA method achieves the best results among these methods. The average improvement of our method is up to $0.24$dB, which is nearly twice that of FR and BN. Visualization results in our Appendix F show that our IGA enables NeRF to capture more accurate and high-frequency reconstruction results, thereby improving PSNR values.
@@ 7 Conclusion
In this paper, we propose an effective gradient adjustment strategy to purposefully improve the spectral bias of multi-layer perceptrons (MLPs) for better implicit neural representations (INRs). We delve into the linear dynamics model of MLPs and theoretically identify that the empirical Neural Tangent Kernel (eNTK) matrix connects spectral bias with the linear dynamics of MLPs. Based on eNTK matrix, we propose our inductive gradient adjustment method via inductive generalization of gradient adjustments from sampled data points. Both theoretical and empirical analysis are conducted to validate impacts of our method on spectral bias. Further, we validate our method on various real-world vision applications of INRs. Our method can effectively tailor improvements on spectral bias and lead to better representation for common INRs network architectures. We hope our study could inspire future works to focus on controlling the bias of neural networks by adjusting training dynamics and yield better performance.
Appendix A Proofs for Section 4
In this section, we give proofs of theorem in the section 4 and more detailed explanation of our theoretical results.
@@ A.1 Analysis framework
where $\nabla_{\Theta}f({\bm{X}};\Theta)\in\mathbb{R}^{q\times N}$; $q$ is the number of parameters. We assume that ${\bm{K}},\tilde{{\bm{K}}}$ are full rank, indicating that they are positive definite. This assumption generally holds due to the complexity of neural networks. Then we have the following standard orthogonal spectral decomposition: ${\bm{K}}=\sum_{i=1}^{N}\lambda_{i}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and $\tilde{{\bm{K}}}=\sum_{i=1}^{N}\tilde{\lambda_{i}}\tilde{{\bm{v}}_{i}}\tilde{{\bm{v}}_{i}}^{\top}$, which eigenvalues are indexed in descending order of magnitude. We construct the corresponding transformation matrices as ${\bm{S}}=\sum_{i=1}^{N}\frac{g_{i}(\lambda_{i})}{\lambda_{i}}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and $\tilde{{\bm{S}}}=\sum_{i=1}^{N}\frac{g_{i}(\tilde{\lambda_{i}})}{\tilde{\lambda_{i}}}\tilde{{\bm{v}}}_{i}\tilde{{\bm{v}}}_{i}^{\top}$. From Jacot et al. (2018); Du et al. (2018); Arora et al. (2019); Geifman et al. (2023), we have that:
@@ A.2 Basic Lemmas
In this section, we show some basic lemmas. Lemma 1 describes the local properties of $f({\bm{x}},\Theta)$. Based on Lemma 1, we derive Lemma 2, 3 and 4, which characterize training dynamics of $f({\bm{x}},\Theta)$ under different adjusted gradients.
Lemma 1.
(Modified from Lemma A.1 of Geifman et al. (2023)) For bounded matrices, i.e., ${\bm{I}},{\bm{S}},\tilde{{\bm{S}}}$, there is a $\kappa>0$ such that for every $C>0$, with high probability over random initialization the following holds: $\forall\Theta,\tilde{\Theta}\in B(\Theta_{0},Cm^{-\frac{1}{2}})$ at time step $t$:
@@
$\|\nabla_{\Theta}f({\bm{X}}){\bm{A}}\|_{F}$ $\leq\frac{\kappa}{\sqrt{m}}$ (15)
@@
As discussed in Lee et al. (2019); Geifman et al. (2023), the core of Lemma 1 is the requirement that the matrix ${\bm{A}}$ is bounded. Therefore, Lemma 1 clearly holds.
Lemma 2.
The parameters are updated by: $\Theta_{t+1}=\Theta_{t}-\eta\nabla_{\Theta_{t}}f({\bm{X}},\Theta_{t}){\bm{S}}{\bm{r}}_{t}$ with $\eta<(min(g(\lambda))+max(g(\lambda)))^{-1}$. For $\epsilon>0$, there always exists $M>0$, when $m>M$, such that with high probability over the random initialization, we have that :
@@
where ${\bm{r}}_{t}=(f({\bm{x}}_{i},\Theta_{t})-y_{i})_{i=1}^{N}$ and $|\xi(t)|<\epsilon$
@@
$\leq^{(1)}\eta LR_{0}\|{\bm{K}}-\tilde{{\bm{K}}}\|_{F}\leq^{(2)}\frac{\epsilon}{2}$
@@
$\leq\frac{\eta\kappa^{2}R_{0}}{m}\|\Theta_{t}-\Theta_{t+1}\|_{2}\leq\frac{\eta\kappa^{2}R_{0}}{m}\|\eta\nabla_{\Theta_{t}}f({\bm{X}},\Theta_{t}){\bm{S}}{\bm{r}}_{t}\|\leq^{(1)}\frac{\eta^{2}\kappa^{3}R_{0}^{2}}{m^{3/2}}$
where $(1)$ follows the Lemma 1. Therefore, we have that ${\bm{r}}_{t+1}=\sum_{i=1}^{N}(1-\eta g(\lambda_{i}))({\bm{v}}_{i}^{\top}{\bm{r}}_{t}){\bm{v}}_{i}\pm\bm{\xi}^{\prime}(t)$; there always exits $M_{1}>0$, such that $m>M_{1}$, $\|\bm{\xi}^{\prime}(t)\|\leq\|A\|+\|B\|<\epsilon$. Further, we have that:
@@
Lemma 3.
The parameters are updated by: $\Theta_{t+1}=\Theta_{t}-\eta\nabla_{\Theta_{t}}f({\bm{X}},\Theta_{t})\tilde{{\bm{S}}}{\bm{r}}_{t}$ with $\eta=(max(g(\tilde{\lambda}))+min(g(\tilde{\lambda})))^{-1}$. For $\epsilon>0$, there always exists $M>0$, when $m>M$, such that with high probability over the random initialization, we have that :
@@
where ${\bm{r}}_{t}=(f({\bm{x}}_{i},\Theta_{t})-y_{i})_{i=1}^{N}$ and $|\tilde{\xi}(t)|<\epsilon$
@@
For $\bm{\tilde{\xi}}^{\prime}(t)$, we follow the same technique in the proof of 2. Then we can have that:
@@
$\leq\frac{\eta\kappa^{2}R_{0}}{m}\|\Theta_{t}-\Theta_{t+1}\|_{2}\leq\frac{\eta^{2}\kappa^{2}R_{0}^{2}}{m^{3/2}}$
@@
$\|{\bm{r}}_{t+1}\|_{2}^{2}$ $=\sum_{i=1}^{N}(1-\eta g(\tilde{\lambda_{i}}))^{2}(\tilde{{\bm{v}}}^{\top}_{i}{\bm{r}}_{t})^{2}+\|\bm{\tilde{\xi}}^{\prime}(t)\|_{2}^{2}\pm 2\bm{\tilde{\xi}}^{\prime}(t)^{\top}\sum_{i=1}^{N}(1-\eta g(\tilde{\lambda_{i}}))(\tilde{{\bm{v}}_{i}}^{\top}{\bm{r}}_{t})\tilde{{\bm{v}}_{i}}$
$=\sum_{i=1}^{N}(1-\eta g(\tilde{\lambda_{i}}))^{2}(\tilde{{\bm{v}}}^{\top}_{i}{\bm{r}}_{t})^{2}+\underbrace{\|\bm{\tilde{\xi}}^{\prime}(t)\|_{2}^{2}\pm 2\bm{\tilde{\xi}}^{\prime}(t)^{\top}({\bm{r}}_{t}\pm\bm{\tilde{\xi}}^{\prime}(t))}_{\tilde{\xi}(t)}$
@@
where $(1),(2)$ follow the Lemma 1 and ${\bm{r}}_{t}$ is bounded by $R_{0}$; Therefore, there exists $M>0$, such that $m>M$, we have that $|\xi^{\prime}_{i}(t)|<\epsilon$
Lemma 2 and lemma 3 show the training dynamics of the residual ${\bm{r}}_{t}$ by NTK-based and eNTK-based gradient adjustments. Equation 16 shows that $\{g_{i}(\lambda_{i})\}_{i=1}^{N}$ control the decay rates of different frequency components of the residual ${\bm{r}}_{t}$ as eigenvectors of ${\bm{K}}$ are the spherical harmonics (Ronen et al., 2019). Despite eNTK-based and NTK-based gradient adjustments have the similar form, they differ in the eigenvector directions $\tilde{{\bm{v}}_{i}}$, which leads to errors. We give a theoretical analysis to this error in 4.1.
@@
(Theorem 4.1 from the paper) The following standard orthogonal spectral decomposition exists: ${\bm{K}}=\sum_{i=1}^{N}\lambda_{i}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and $\tilde{{\bm{K}}}=\sum_{i=1}^{N}\tilde{\lambda_{i}}\tilde{{\bm{v}}_{i}}\tilde{{\bm{v}}_{i}}^{\top}$, satisfying that ${\bm{v}}_{i}^{\top}\tilde{{\bm{v}}}_{i}>0$ for $i\in[N]$. We denoted as $max_{i}\{min\{\lambda_{i}-\lambda_{i+1},\lambda_{i+1}-\lambda_{i+2}\}\}=G$. Let $\{g_{i}(x)\}_{i=1}^{N}$ be a set of Lipschitz continuous functions, with the Supremum of their Lipschitz constants denoted by $k$, with $\eta<min\{(max(g(\lambda))+min(g(\lambda)))^{-1},(max(g(\tilde{\lambda}))+min(g(\tilde{\lambda})))^{-1}\}$, for $\epsilon>0$, there always exists $M>0$, such that $m>M$, for $i\in[N]$, we have that $|g_{i}(\lambda_{i})-g_{i}(\tilde{\lambda_{i}})|<\epsilon_{1},\|{\bm{v}}_{i}-\tilde{{\bm{v}}}_{i}\|<\epsilon_{2}$; furthermore, we have that:
@@
$\|{\bm{E}}\|_{2}=\sqrt{(\lambda^{\prime}_{1})^{2}}\leq\sqrt{\sum_{i=1}^{N}(\lambda^{\prime}_{i})^{2}}=\sqrt{trace({\bm{E}}^{\top}{\bm{E}})}=\|{\bm{E}}\|_{F}$
@@
$|g_{i}(\lambda_{i})-g_{i}(\tilde{\lambda}_{i})|\leq k|\lambda_{i}-\tilde{\lambda}_{i}|<k\epsilon$
@@
$\|\tilde{{\bm{v}}}_{i}-{\bm{v}}_{i}\|\leq\frac{2^{3/2}\|{\bm{E}}\|_{2}}{G}<\frac{2^{3/2}\epsilon}{G},i\in[N]$
@@
$<|1-\eta g_{i}(\lambda_{i})|\cdot(2\|{\bm{v}}_{i}^{\top}\|+\|\tilde{{\bm{v}}}_{i}^{\top}-{\bm{v}}_{i}\|_{2})\|\tilde{{\bm{v}}}_{i}^{\top}-{\bm{v}}_{i}\|_{2}\|{\bm{r}}_{t}\|_{2}^{2}<\epsilon\frac{8R_{0}^{2}(G\|{\bm{v}}_{i}^{\top}\|+\epsilon)}{G^{2}}$
@@
$B<\eta|g_{i}(\tilde{\lambda}_{i})-g_{i}(\lambda_{i})|(\tilde{{\bm{v}}}_{i}^{\top}{\bm{r}}_{t})^{2}<\epsilon k\eta(\|\tilde{{\bm{v}}}_{i}^{\top}-{\bm{v}}_{i}^{\top}\|\|{\bm{r}}_{t}\|+\|{\bm{v}}_{i}\|\|{\bm{r}}_{t}\|)^{2}<\epsilon k\eta R_{0}(\frac{2^{3/2}}{G}\epsilon+v)^{2}$
@@
$|(1-\eta g_{i}(\lambda_{i}))^{2}({\bm{v}}_{i}^{\top}{\bm{r}}_{t})^{2}-(1-\eta g_{i}(\tilde{\lambda_{i}}))^{2}(\tilde{{\bm{v}}_{i}}^{\top}{\bm{r}}_{t})^{2}|<\frac{8R_{0}}{G^{2}}(k\eta\epsilon^{3}+(v+1)\epsilon^{2}+(v^{2}+Gv)\epsilon)$
@@
Combing this result with the Lemma 2 and 3, there always exists $M>max\{M_{0},M_{1},M_{2}\}$, we have that:
@@
$<\epsilon N((\frac{16vR_{0}}{G}+\eta kv^{2}R_{0})+\epsilon^{2})+2\epsilon$
@@
(Theorem 4.2 from the paper) ${\bm{X}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$ are partitioned by order into $n$ groups, where each group is ${\bm{X}}_{j}=\{{\bm{x}}^{j}_{1},\dots,{\bm{x}}^{j}_{p}\}$ and $N=np$. For each group, one data point is sampled denoted as ${\bm{x}}^{j}$ and these $n$ data points form ${\bm{X}}_{e}=\{{\bm{x}}^{j}\}_{j=1}^{n},n\ll N$. There exists $\epsilon>0$, for $j=1,\dots,n$ and any $1\leq i_{1},i_{2}\leq p$, such that $|(f({\bm{x}}^{j}_{i_{1}},\Theta_{t})-y^{j}_{i_{1}})-(f({\bm{x}}^{j}_{i_{2}},\Theta_{t})-y^{j}_{i_{2}})|,\|\nabla_{\Theta_{t}}f({\bm{x}}^{j}_{i_{1}})-\nabla_{\Theta_{t}}f({\bm{x}}^{j}_{i_{2}})\|<\epsilon$. Then, for any ${\bm{x}}_{i}$, assumed that ${\bm{x}}_{i}$ belongs to ${\bm{X}}_{j_{i}}$, such that:

### v2

@@ 5.3 Learning 5D neural radiance fields
Learning neural radiance fields for novel view synthesis, i.e., NeRF, is the main application of INRs (Mildenhall et al., 2021; Saragadam et al., 2023; Xie et al., 2023; Shi et al., 2024a). The main process of NeRF is to build a neural radiance field from a 5D coordinate space to RGB space.
Specifically, given a ray $i$ from the camera into the scene, the MLP takes the 5D coordinates (3D spatial locations and 2D view directions) of points along the ray as inputs and outputs the corresponding color $\mathbf{c}$ and volume density $\sigma$. Then $\mathbf{c}$ and $\sigma$ are combined using numerical volume rendering to obtain the final pixel color of the ray $i$. Following our sampling strategy, We place the 2D view directions last during flattening and set $p$ as the number of sampling points per ray. That is, each ray is treated as a group. For each group, we sample the point with the maximum integral weight. We apply our method to the original NeRF (Mildenhall et al., 2021). The “NeRF-PyTorch” codebase (Yen-Chen, 2020) is used, and we follow its default settings for all methods. More details can be found in Appendix I.
Table 3: Average metrics of 5D neural radiance fields by different methods. Per-scene results are provided in Table. 14.
@@
Table 3 lists average metrics of four methods on the down-scaled Blender dataset (Mildenhall et al., 2021). Our IGA method achieves the best results among these methods. The average improvement of our method is up to $0.24$dB, which is nearly twice that of FR and BN. Visualization results in our Appendix I show that our IGA enables NeRF to capture more accurate and high-frequency reconstruction results.
5.4 Time and memory analysis
In this subsection, we analyze the training time and memory cost of our IGA in the aforementioned experiments. For both vanilla gradient descent and IGA, we present the average per-iteration training time and memory cost over all models within each experiment, as summarized in Table 4. We only report the results of adjustments using the complete eNTK matrix (+Full eNTK) for 1D simple function approximation, due to its excessive memory cost (exceeding $80$GiB) and prohibitive training time in other INR tasks.
Table 4: Average per-iteration training time and memory cost by different methods. “Vanilla” denotes standard training using conventional gradients; “+Full eNTK” denotes training with gradient adjustment based on the complete eNTK matrix ${\bm{K}}$; “+IGA” denotes training with inductive gradient adjustment. The abbreviations “1D”, “2D”, “3D”, and “5D” are used to represent the experiments on simple functions, 2D images, 3D shapes, and neural radiance fields, respectively.
INR Training time (s) Memory cost (MB)
tasks Vanilla +Full eNTK +IGA Vanilla +Full eNTK +IGA
1D 0.028 0.185 0.044 965 5861 1574
2D 0.061 00 0.088 3703 00 5174
3D 0.118 00 0.169 2291 00 2468
5D 0.159 00 0.286 9451 00 11033
Overall, IGA averages 1.56×1.56$\times$1.56 × the training time and 1.34×1.34$\times$1.34 × the memory of the baseline but achieves greater improvements (on average, 2.0×2.0$\times$2.0 × those of prior FR and BN). Compared to “+Full eNTK”, IGA reduces the training time and memory consumption by at least 4×4$\times$4 ×.
@@ 6 Conclusion
In this paper, we propose an effective gradient adjustment strategy to purposefully improve the spectral bias of multi-layer perceptrons (MLPs) for better implicit neural representations (INRs). We delve into the linear dynamics model of MLPs and theoretically identify that the empirical Neural Tangent Kernel (eNTK) matrix connects spectral bias with the linear dynamics of MLPs. Based on eNTK matrix, we propose our inductive gradient adjustment method via inductive generalization of gradient adjustments from sampled data. Both theoretical and empirical analysis are conducted to validate impacts of our method on spectral bias. Further, we validate our method on various real-world vision applications of INRs. Our method can effectively tailor improvements on spectral bias and lead to better representation for common INRs network architectures. We hope this study inspires further work on neural network training dynamics to improve application performance.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
This work was supported by National Natural Science Foundation of China (No. 62476051) and Sichuan Natural Science Foundation (No. 2024NSFTD0041).
Appendix
This appendix is organized as follows. In Sec. A, we provide the formal version and detailed proof of Theorem 3.1 and 3.2 in the main paper. Then, more empirical evidence from Sec. 4 are provided in Sec. B to support theoretical results in Sec. 3.3. In Sec. C, the visual illustration of our sampling strategy is provided. In Sec. D, a further discussion about our inductive gradient adjustment (IGA) with Adam optimizer referred in Sec. 3.4 is presented. Then, we add experiments of IGA on recent state-of-the-art activations in Sec. E. In Sec. F, the ablation study on sampling strategy is provided. Lastly, we give detailed training settings and per-scene results of our three vision applications in Sec. G, H, I, respectively.
Appendix A Theoretical Analysis
In this section, we first introduce the detailed framework for theoretical analysis. Then, we provide the formal statements and proofs of Theorem 3.1 and 3.2 of Sec. 3.3.
@@ A.1 Analysis framework
where $\nabla_{\Theta}f({\bm{X}};\Theta)\in\mathbb{R}^{q\times N}$; $q$ is the number of parameters. We assume that ${\bm{K}},\tilde{{\bm{K}}}$ are full rank, indicating that they are positive definite. This assumption generally holds due to the complexity of neural networks. Then we have the following standard orthogonal spectral decomposition: ${\bm{K}}=\sum_{i=1}^{N}\lambda_{i}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and $\tilde{{\bm{K}}}=\sum_{i=1}^{N}\tilde{\lambda_{i}}\tilde{{\bm{v}}_{i}}\tilde{{\bm{v}}_{i}}^{\top}$, which eigenvalues are indexed in descending order of magnitude. We construct the corresponding transformation matrices as ${\bm{S}}=\sum_{i=1}^{N}\frac{g_{i}(\lambda_{i})}{\lambda_{i}}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and $\tilde{{\bm{S}}}=\sum_{i=1}^{N}\frac{g_{i}(\tilde{\lambda_{i}})}{\tilde{\lambda_{i}}}\tilde{{\bm{v}}}_{i}\tilde{{\bm{v}}}_{i}^{\top}$ through a set of Lipschitz continuous functions $\{g_{i}(\lambda)\}_{i=1}^{N}$ From Jacot et al. (2018); Du et al. (2018); Arora et al. (2019); Geifman et al. (2023), we have that:
@@ A.2 Basic Lemmas
In this section, we show some basic lemmas. Lemma A.1 describes the local properties of $f({\bm{x}},\Theta)$. Based on Lemma A.1, we derive Lemma A.2, A.3 and A.4, which characterize training dynamics of $f({\bm{x}},\Theta)$ under different adjusted gradients.
Lemma A.1.
(Modified from Lemma A.1 of (Geifman et al., 2023)) For bounded matrices, i.e., ${\bm{I}},{\bm{S}},\tilde{{\bm{S}}}$, there is a $\kappa>0$ such that for every $C>0$, with high probability over random initialization the following holds: $\forall\Theta,\tilde{\Theta}\in B(\Theta_{0},Cm^{-\frac{1}{2}})$ at time step $t$:
@@
$\|\nabla_{\Theta}f({\bm{X}}){\bm{A}}\|_{F}$ $\leq\frac{\kappa}{\sqrt{m}},$ (14)
@@
As discussed in Lee et al. (2019); Geifman et al. (2023), the core of Lemma A.1 is the requirement that the matrix ${\bm{A}}$ is bounded. Therefore, Lemma A.1 clearly holds. Then, to better illustrate our core theoretical analysis, we present the following Lemma A.2, A.3 and A.4. Please note that a similar version of Lemma A.2 has been proved in (Geifman et al., 2023). However, it focuses on ${\bm{r}}_{t+1}$ induced by only ${\bm{S}}$ and the details of the eigenvalues between ${\bm{r}}_{t+1}$ and ${\bm{r}}_{t}$ are omitted. Therefore, our Lemma A.2, A.3 and A.4 characterize the dynamics in one iteration from ${\bm{r}}_{t}$ to ${\bm{r}}_{t+1}$ and eigenvalues induced by NTK-based and eNTK-based adjustments and vanilla gradient, respectively.
Lemma A.2.
The parameters are updated by: $\Theta_{t+1}=\Theta_{t}-\eta\nabla_{\Theta_{t}}f({\bm{X}},\Theta_{t}){\bm{S}}{\bm{r}}_{t}$ with $\eta<(min(g_{i}(\lambda))+max(g_{i}(\lambda)))^{-1}$. For $\epsilon>0$, there always exists $M>0$, when $m>M$, such that with high probability over the random initialization, we have that :
@@
where ${\bm{r}}_{t}=(f({\bm{x}}_{i},\Theta_{t})-y_{i})_{i=1}^{N}$ and $|\xi(t)|<\epsilon$.
@@
$\leq^{(1)}\eta LR_{0}\|{\bm{K}}-\tilde{{\bm{K}}}\|_{F}\leq^{(2)}\frac{\epsilon}{2},$
@@
$\leq\frac{\eta\kappa^{2}R_{0}}{m}\|\Theta_{t}-\Theta_{t+1}\|_{2}\leq\frac{\eta\kappa^{2}R_{0}}{m}\|\eta\nabla_{\Theta_{t}}f({\bm{X}},\Theta_{t}){\bm{S}}{\bm{r}}_{t}\|\leq^{(1)}\frac{\eta^{2}\kappa^{3}R_{0}^{2}}{m^{3/2}},$
where $(1)$ follows the Lemma A.1. Therefore, we have that ${\bm{r}}_{t+1}=\sum_{i=1}^{N}(1-\eta g_{i}(\lambda_{i}))({\bm{v}}_{i}^{\top}{\bm{r}}_{t}){\bm{v}}_{i}\pm\bm{\xi}^{\prime}(t)$; there always exits $M_{1}>0$, such that $m>M_{1}$, $\|\bm{\xi}^{\prime}(t)\|\leq\|A\|+\|B\|<\epsilon$. Further, we have that:
@@
Lemma A.3.
The parameters are updated by: $\Theta_{t+1}=\Theta_{t}-\eta\nabla_{\Theta_{t}}f({\bm{X}},\Theta_{t})\tilde{{\bm{S}}}{\bm{r}}_{t}$ with $\eta=(max(g_{i}(\tilde{\lambda}))+min(g_{i}(\tilde{\lambda})))^{-1}$. For $\epsilon>0$, there always exists $M>0$, when $m>M$, such that with high probability over the random initialization, we have that :
@@
where ${\bm{r}}_{t}=(f({\bm{x}}_{i},\Theta_{t})-y_{i})_{i=1}^{N}$ and $|\tilde{\xi}(t)|<\epsilon$.
@@
For $\bm{\tilde{\xi}}^{\prime}(t)$, we follow the same technique in the proof of A.2. Then we can have that:
@@
$\leq\frac{\eta\kappa^{2}R_{0}}{m}\|\Theta_{t}-\Theta_{t+1}\|_{2}\leq\frac{\eta^{2}\kappa^{2}R_{0}^{2}}{m^{3/2}}.$
@@
$\|{\bm{r}}_{t+1}\|_{2}^{2}$ $=\sum_{i=1}^{N}(1-\eta g_{i}(\tilde{\lambda_{i}}))^{2}(\tilde{{\bm{v}}}^{\top}_{i}{\bm{r}}_{t})^{2}+\|\bm{\tilde{\xi}}^{\prime}(t)\|_{2}^{2}\pm 2\bm{\tilde{\xi}}^{\prime}(t)^{\top}\sum_{i=1}^{N}(1-\eta g_{i}(\tilde{\lambda_{i}}))(\tilde{{\bm{v}}_{i}}^{\top}{\bm{r}}_{t})\tilde{{\bm{v}}_{i}}$
$=\sum_{i=1}^{N}(1-\eta g_{i}(\tilde{\lambda_{i}}))^{2}(\tilde{{\bm{v}}}^{\top}_{i}{\bm{r}}_{t})^{2}+\underbrace{\|\bm{\tilde{\xi}}^{\prime}(t)\|_{2}^{2}\pm 2\bm{\tilde{\xi}}^{\prime}(t)^{\top}({\bm{r}}_{t}\pm\bm{\tilde{\xi}}^{\prime}(t))}_{\tilde{\xi}(t)}.$
@@
where $(1),(2)$ follow the Lemma A.1 and ${\bm{r}}_{t}$ is bounded by $R_{0}$; Therefore, there exists $M>0$, such that $m>M$, we have that $|\xi^{\prime}_{i}(t)|<\epsilon$. ∎
Lemma A.2 and lemma A.3 show the training dynamics of the residual ${\bm{r}}_{t}$ by NTK-based and eNTK-based gradient adjustments. Equation 15 shows that $\{g_{i}(\lambda_{i})\}_{i=1}^{N}$ control the decay rates of different frequency components of the residual ${\bm{r}}_{t}$ as eigenvectors of ${\bm{K}}$ are the spherical harmonics (Ronen et al., 2019). Lemma A.4 show the training dynamics of residual at data point ${\bm{x}}_{i}$ by vanilla gradient descent. Despite eNTK-based and NTK-based gradient adjustments have the similar form, they differ in the eigenvector directions $\tilde{{\bm{v}}_{i}}$, which leads to errors. We give a theoretical analysis to this error in Theorem 3.1.
@@
(formal version of Theorem 3.1 from the paper) The following standard orthogonal spectral decomposition exists: ${\bm{K}}=\sum_{i=1}^{N}\lambda_{i}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and $\tilde{{\bm{K}}}=\sum_{i=1}^{N}\tilde{\lambda_{i}}\tilde{{\bm{v}}_{i}}\tilde{{\bm{v}}_{i}}^{\top}$, satisfying that ${\bm{v}}_{i}^{\top}\tilde{{\bm{v}}}_{i}>0$ for $i\in[N]$. We denoted as $\max_{i}\{\min\{\lambda_{i}-\lambda_{i+1},\lambda_{i+1}-\lambda_{i+2}\}\}=G$. Let $\{g_{i}(x)\}_{i=1}^{N}$ be a set of Lipschitz continuous functions, with the Supremum of their Lipschitz constants denoted by $k$, with $\eta<\min\{(\max(g_{i}(\lambda))+\min(g_{i}(\lambda)))^{-1},(\max(g_{i}(\tilde{\lambda}))+\min(g_{i}(\tilde{\lambda})))^{-1}\}$, for $\epsilon>0$, there always exists $M>0$, such that $m>M$, for $i\in[N]$, we have that $|g_{i}(\lambda_{i})-g_{i}(\tilde{\lambda_{i}})|<\epsilon_{1},\|{\bm{v}}_{i}-\tilde{{\bm{v}}}_{i}\|<\epsilon_{2}$; furthermore, we have that:
@@
$\|{\bm{E}}\|_{2}=\sqrt{(\lambda^{\prime}_{1})^{2}}\leq\sqrt{\sum_{i=1}^{N}(\lambda^{\prime}_{i})^{2}}=\sqrt{trace({\bm{E}}^{\top}{\bm{E}})}=\|{\bm{E}}\|_{F},$
@@
$|g_{i}(\lambda_{i})-g_{i}(\tilde{\lambda}_{i})|\leq k|\lambda_{i}-\tilde{\lambda}_{i}|<k\epsilon.$
@@
$\|\tilde{{\bm{v}}}_{i}-{\bm{v}}_{i}\|\leq\frac{2^{3/2}\|{\bm{E}}\|_{2}}{G}<\frac{2^{3/2}\epsilon}{G},i\in[N].$
@@
$<|1-\eta g_{i}(\lambda_{i})|\cdot(2\|{\bm{v}}_{i}^{\top}\|+\|\tilde{{\bm{v}}}_{i}^{\top}-{\bm{v}}_{i}\|_{2})\|\tilde{{\bm{v}}}_{i}^{\top}-{\bm{v}}_{i}\|_{2}\|{\bm{r}}_{t}\|_{2}^{2}<\epsilon\frac{8R_{0}^{2}(G\|{\bm{v}}_{i}^{\top}\|+\epsilon)}{G^{2}}.$
@@
$B<\eta|g_{i}(\tilde{\lambda}_{i})-g_{i}(\lambda_{i})|(\tilde{{\bm{v}}}_{i}^{\top}{\bm{r}}_{t})^{2}<\epsilon k\eta(\|\tilde{{\bm{v}}}_{i}^{\top}-{\bm{v}}_{i}^{\top}\|\|{\bm{r}}_{t}\|+\|{\bm{v}}_{i}\|\|{\bm{r}}_{t}\|)^{2}<\epsilon k\eta R_{0}(\frac{2^{3/2}}{G}\epsilon+v)^{2}.$
@@
$|(1-\eta g_{i}(\lambda_{i}))^{2}({\bm{v}}_{i}^{\top}{\bm{r}}_{t})^{2}-(1-\eta g_{i}(\tilde{\lambda_{i}}))^{2}(\tilde{{\bm{v}}_{i}}^{\top}{\bm{r}}_{t})^{2}|<\frac{8R_{0}}{G^{2}}(k\eta\epsilon^{3}+(v+1)\epsilon^{2}+(v^{2}+Gv)\epsilon).$
@@
Combing this result with the Lemma A.2 and A.3, there always exists $M>max\{M_{0},M_{1},M_{2}\}$, we have that:
@@
$<\epsilon N((\frac{16vR_{0}}{G}+\eta kv^{2}R_{0})+\epsilon^{2})+2\epsilon.$
@@
(formal version of Theorem 3.2 from the paper) ${\bm{X}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$ is partitioned into $n$ groups, where each group is ${\bm{X}}_{j}=\{{\bm{x}}^{j}_{1},\dots,{\bm{x}}^{j}_{p}\}$ and $N=np$. For each group, one data point is sampled denoted as ${\bm{x}}^{j}$ and these $n$ data points form ${\bm{X}}_{e}=\{{\bm{x}}^{j}\}_{j=1}^{n},n\ll N$. There exists $\epsilon>0$, for $j=1,\dots,n$ and any $1\leq i_{1},i_{2}\leq p$, such that $|(f({\bm{x}}^{j}_{i_{1}},\Theta_{t})-y^{j}_{i_{1}})-(f({\bm{x}}^{j}_{i_{2}},\Theta_{t})-y^{j}_{i_{2}})|,\|\nabla_{\Theta_{t}}f({\bm{x}}^{j}_{i_{1}})-\nabla_{\Theta_{t}}f({\bm{x}}^{j}_{i_{2}})\|<\epsilon$. Then, for any ${\bm{x}}_{i}$, assumed that ${\bm{x}}_{i}$ belongs to ${\bm{X}}_{j_{i}}$, such that: